# Optimas: Optimizing Compound AI Systems with Globally Aligned Local Rewards

### v3

@@ Appendix B Theoretical Analysis
$\arg\min_{p}\mathbb{E}[\log(\sigma_{1}(y\cdot p(x)))]=p^{*}.$
@@ Appendix D Compound AI System Details
Amazon Session Analyzer Qwen 2.5 1B Model Params PPO (RL)
Candidate Profiler Qwen 2.5 1B Model Params PPO (RL)
@@
Session Analyzer and Candidate Profiler both use the Qwen 2.5 1B model; we optimize their model parameters with PPO reinforcement learning (Schulman et al., 2017). This helps each module better encode task-specific knowledge, i.e., user sessions and product candidates. The final Next Item Decider is a GPT-4o-mini module, whose prompt we optimize.
@@
Four GPT-4o-mini modules (Question Rewriter, Info Extractor, Hint Generator, Answer Generator) each rely on prompt optimization to improve multi-hop reasoning. Meanwhile, the Retriever is retriever with a key hyperparameter k, the number of passages to pull. We search $\texttt{k}\in\{1,5,10,25\}$ which we also tune via global hyperparameter search across training instances.
@@ Appendix E Baseline Details
  • LLMSelector (Chen et al., 2025b): A lightweight policy selects the best LLM per component via model routing, without updating other configurations. Only applicable on PubMedQA.
    We run LLMSelector (Chen et al., 2025b) with the LLMDiagnoser to estimate per-module performance, following their procedure. We perform two rounds of allocation updates with 100 training examples each round.
  • Hierarchical Behavior Cloning (HBC) (Le et al., 2018): A hierarchical imitation learning method that optimizes components to produce outputs similar to those that lead to high global rewards.
@@ Appendix G More Experiment Results
G.1 Optimas Data efficiency
Figure 7: Impact of training data size on evaluation accuracy for the HotpotQA system. Optimas maintains robust performance even with reduced training data, demonstrating only a 5.19% absolute performance drop when using just 12.5% of the data compared to the full dataset.
To evaluate the data efficiency of Optimas, we examine how the amount of training data (system runs) affects the performance of learned LRFs. We train local reward models on the HotpotQA system using varying percentages of available training data: 12.5%, 25%, 50%, 75%, and 100%. Each model is trained for 3 epochs, and performance is measured by pairwise ranking accuracy—how often an output with higher global reward is assigned a higher local reward compared to an output with lower global reward. Optimas maintains reasonable performance across different data sizes as show in Figure 7. Even with only 12.5% of the data, the system achieves 65.46% accuracy (92.7% of full-data performance), while 25% of the data yields 67.22% accuracy (95.1% of full-data performance). The diminishing returns beyond 25% data suggest that the core patterns for local-global alignment can be captured efficiently from limited preference data.
This data efficiency is particularly valuable in practice, where collecting extensive preference data through system runs can be computationally expensive. Combined with the reduced computational cost of local optimization, these results demonstrate that Optimas provides a practical solution for optimizing complex compound AI systems even in resource-constrained environments.
G.2 Local vs. Global reward landscapes
@@ G.1 Optimas Data efficiency
G.3 LRF backbone size
Table 9: On HotpotQA, effect of LRF backbone size on pairwise ranking accuracy (%). Lightweight backbones are competitive with larger ones.
Backbone size Pairwise accuracy (%)
1B 72.58
3B 70.04
8B 71.20
We also examine whether the size of the LRF backbone affects alignment quality. On HotpotQA, we compare 1B, 3B, and 8B backbones (Table 9). Performance remains comparable across sizes, with pairwise accuracies of 72.58%, 70.04%, and 71.20%, respectively. The lack of clear benefit from larger backbones suggests that lightweight models are sufficient for this task, making local optimization both effective and computationally efficient.
G.4 Prompt candidates per step

### v4

@@ Appendix B Theoretical Analysis
$\arg{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{max}}_{p}\mathbb{E}[\log(\sigma_{1}(y\cdot p(x)))]=p^{*}.$
@@ Appendix D Compound AI System Details
Amazon Session Analyzer Qwen 2.5 1.5B Model Params PPO (RL)
Candidate Profiler Qwen 2.5 1.5B Model Params PPO (RL)
@@
Session Analyzer and Candidate Profiler both use the Qwen 2.5 1.5B model; we optimize their model parameters with PPO reinforcement learning (Schulman et al., 2017). This helps each module better encode task-specific knowledge, i.e., user sessions and product candidates. The final Next Item Decider is a GPT-4o-mini module, whose prompt we optimize.
@@
Four GPT-4o-mini modules (Question Rewriter, Info Extractor, Hint Generator, Answer Generator) each rely on prompt optimization to improve multi-hop reasoning. Meanwhile, the Retriever is a retriever with a key hyperparameter k, the number of passages to pull. We search $\texttt{k}\in\{1,5,10,25\}$ which we also tune via global hyperparameter search across training instances.
@@ Appendix E Baseline Details
  • LLMSelector (Chen et al., 2025b): A lightweight policy selects the best LLM per component via model routing, without updating other configurations. Only applicable on PubMedQA. We run LLMSelector (Chen et al., 2025b) with the LLMDiagnoser to estimate per-module performance, following their procedure. We perform two rounds of allocation updates with 100 training examples each round.
  • REINFORCE (Williams, 1992): A policy-gradient baseline that directly updates the parameters of local LLM components to maximize the task reward. This method is only applicable on Amazon, where we deploy two locally hosted LLMs with trainable parameters. We optimize each component with REINFORCE using sampled trajectories from the full system and propagate the scalar reward back to the corresponding module-level policy. We use 16 rollouts per step to estimate the expected system performance, with learning rate $10^{-5}$ for 100 training steps.
  • Hierarchical Behavior Cloning (HBC) (Le et al., 2018): A hierarchical imitation learning method that optimizes components to produce outputs similar to those that lead to high global rewards. Specifically, we collect successful trajectories to approximate ground truth intermediate outputs, and then perform supervised updates on the local components to mimic/clone the ideal behavior.
@@ Appendix G More Experiment Results
G.1 Local vs. Global reward landscapes
@@
G.2 Prompt candidates per step
@@ G.3 New inputs for LRF adaptation
G.4 Cost Estimation
We report a detailed breakdown of the cost on the Amazon system. This is the only system where components have trainable local models and thus requires PPO training. For all methods, we measure cost in terms of full system runs, i.e., one invocation of the system and assume that all components contribute equally to the per-run cost. We label the three components on Amazon system as $A,B,C$ in topological order.
Table 11: Cost–performance comparison on the Amazon system. Individual cost components are shown in actual runs, with total runs measured in thousands (k). “LRF training” and “LRF adaptation” only apply to Optimas, which learns trainable local reward functions. DSPy and TextGrad optimize only the prompt of the last component $C$, and we report their cost as an effective number of full system runs (see text).
Method LRF training LRF adaptation Validation cost Total runs (k) Performance (%)
Optimas 120 100 93 0.31 24.24
TextGrad 320 0.32 20.88
DSPy 240 0.24 18.18
For Optimas, the total of $\approx 0.31$k effective full system runs decomposes into three parts: (i) LRF training: we collect 60 initial preference pairs per component; each pair requires 2 full system runs, giving $60\times 2=120$ runs; (ii) LRF adaptation: during local optimization, we update the LRFs 5 times, each time collecting 10 new preference pairs per component, again with 2 runs per pair, for a total of $5\times 10\times 2=100$ runs; and (iii) global validation: whenever a local update is predicted to improve the global metric, we evaluate the new configuration on a held-out validation set. We use 20 validation inputs and observe 7 such updates, but the effective cost is discounted by a factor of $2/3$ due to cached trajectories:
$20\times 7\times\tfrac{2}{3}\approx 93.$
The discount factor arises because, for each validation, we do not always need to downstream sample the entire system: roughly speaking, if $A$ is updated, we need 1 system run per validation input; if $B$ is updated, we need $2/3$ system run per validation input; if $C$ is updated, we need $1/3$ system run per validation input
$\frac{1+\tfrac{2}{3}+\tfrac{1}{3}}{3}=\tfrac{2}{3}$
of a full system run. The resulting global performance is $24.24\%$.
For DSPy and TextGrad, we follow their standard setup and optimize only the prompt for the final component $C$, while holding $A$ and $B$ fixed. To obtain a fair cost comparison in terms of effective full system runs, we (i) pre-compute a pool of 20 validation inputs by running $A$ and $B$ once, and then (ii) run only $C$ during optimization. Assuming $A$, $B$, and $C$ have comparable cost, we convert the number of $C$-only calls into an effective number of full system runs by dividing by three. With 48 optimization steps for TextGrad and 36 for DSPy, this yields
$20\times 48/3=320\quad\text{and}\quad 20\times 36/3=240$
effective full system runs, corresponding to 0.32k and 0.24k in Table 11. These are conservative lower bounds, since any additional evaluation of upstream components would only increase their cost.
Under this cost-normalized view, Optimas achieves the best global performance while using a comparable (slightly lower) effective number of full system runs than TextGrad and only moderately more than DSPy. This demonstrates that local optimization with learned reward functions can deliver stronger performance without incurring prohibitive cost.
Additional PPO training cost.
Beyond system-run cost, Optimas incurs additional compute for training local models with PPO. We measure this cost in GPU-hours. For each local PPO update on the Amazon system, we train for 3 epochs on a single NVIDIA A100-SXM4-80GB GPU. Averaged over 5 runs of local optimization, each run takes approximately 12 minutes, yielding a total of about 6 GPU-hours (equivalently, 1.5 hours on 4 GPUs). This one-time training overhead is modest relative to the cost of repeated system evaluations and is only required for systems with trainable local components.